INAF queue observing, 2013 June 14-15

Observer: F. Cusano, M. Fumana
Telescope Operator: G. Becchetti
Support Astronomer: O. Kuhn (on call)
Instrument: MODS


At the sunset heavy clouds. Sky almost covered, in the North direction only thin cirrus.
Around 3.30UT sky looks better. Slight chance of storm.
We open the dome an hour later because of the whether conditions.
Until 8.30UT exposures are good. Around ~8.50 we had a problem of communication and we were not guiding. We solved the problem, but clouds came.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Sommariva_31 1625-2648 Blue grating 3.5h 0011 exposure no guiding
Standard bd+24 Dual grating    
Israel Cats191043 Dual grating ~0.3h Sky brightness raised


Start using MODS

04:55 Dome opened
  • Starting collimation

05:04 acqMODS n1625.acq

05:17 execMODS n1625_13.obs 0005-0006
  • seeing ~0.9"
  • Sky is almost clear now

06:22 execMODS n1625_14.obs 0007-0008

07:26 execMODS n1625_15.obs 0009-0010
  • seeing <1"
  • seeing ~1.2"

08:31 execMODS n1625_16.obs 0011
  • seeing 1.2
  • WFS problem. 1st exposure stopped after 1680s
  • In the second part (~10min) of this exposure we are tracking, but we are not guiding

09:02 Clear active optics
  • send preset again . acqMODS n1625.acq. Failed
  • Communication problem
  • Reset MODS agw. Restored communication

09:14 acqMODS n1625.acq

09:23 execMODS n1625_16.obs 0012
  • clouds are coming. Still guiding

*09:57 Switch to Israel. Standard
  • acqMODS bd+24.acq
  • Adjust offset by hand Dx=-1.59, Dy=-1.0, exptime=3s

10:10 execMODS bd+24.obs

*10:20*acqMODS cats191043.acq

10:30 execMODS cats191043.obs blue 0021, red 0013
  • Stop 1st exposure after 1117s (sky brightness is raising)

11:10 Dome closed

-- MarcoFumana - 15 Jun 2013
Topic revision: r4 - 25 Sep 2013, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback